OpenStudy (greengalaxy):

Complete the statement. -2(3+6)= x - 12 A. –6 B. –3 C. 3 D. 6

OpenStudy (greengalaxy):

@rational

OpenStudy (kidrah69):

Your first step is to distribute.

OpenStudy (kidrah69):

\[\huge\rm-2(3+6)\]

OpenStudy (greengalaxy):

Oh my gosh idk how I forgot that.. okay so -2 x 3?

OpenStudy (anonymous):

x=−6

OpenStudy (kidrah69):

@GreenGalaxy yep

OpenStudy (greengalaxy):

Thank you guys I get it now -2 + 3 -2 + 6 Right

OpenStudy (kidrah69):

Simplify 3+6 to 9 −2×9=x−12 2 Simplify 2×9 to 18 −18=x−12 3 Add 12 to both sides −18+12=x 4 Simplify −18+12 to −6 −6=x 5 Switch sides x=−6

OpenStudy (kidrah69):

Do you understand now? :)

OpenStudy (greengalaxy):

Oh snap! I didn't remeber that! Thanks! I gave you a medal.

OpenStudy (kidrah69):

haha thanks :3 and no probs

OpenStudy (greengalaxy):

Going to close this question